1.4263mV to kV
2. 2.95km² to mm²
1
Expert's answer
2020-08-13T12:25:05-0400

1. First, let's convert 4263mV to volts ("1mV = 10^{-3}V" ):

"4263\space mV = 4263\times 10^{-3}V"

Then convert to kV ("1V = 10^{-3}kV" ):

"4263\times 10^{-3}V =4263\times 10^{-6}kV = 4.236\times 10^{-3}kV"

2. First, let's convert 2.95km² to m^2 ("1km = 10^{6}m^2" ):

"2.95km² = 2.95\times 10^{6}m^2"

Then convert to mm^2 ("1m^2 = 10^{6}mm^2" ):

"2.95\times 10^{6}m^2 =2.95\times 10^{6}\times10^6 mm^2 = 2.95\times 10^{12}mm^2"

Answer. 1. "4.236\times 10^{-3}kV", 2. "2.95\times 10^{12}mm^2".